OpenStudy (anonymous):

write a rule for the following function x -1,0,1,2 y 3,1,1/3,1/9

OpenStudy (kc_kennylau):

1. Realise that y are all powers of 3 2. Realise that y and x are going the reverse direction, so the index must contain -x

OpenStudy (kc_kennylau):

Notice that \(\begin{matrix}x&-1&0&1&2\\\mbox{index of y}&1&0&-1&-2\end{matrix}\)

OpenStudy (kc_kennylau):

And notice that \(\begin{matrix}x&-1&0&1&2\\-x&1&0&-1&-2\end{matrix}\)

OpenStudy (kc_kennylau):

\[\Large\begin{array}{c|c|c|c|c}x&&-1&0&1&2\\\hline-x&&1&0&-1&-2\\\hline3^{-x}&&3&0&\frac13&\frac19\end{array}\]

OpenStudy (kc_kennylau):

sorry the second number of the row \(\large3^{-x}\) should be 1

OpenStudy (anonymous):

okay i think i understand now thank you very much

OpenStudy (kc_kennylau):

no problem :)